OpenStudy (anonymous):

Write the slope-intercept equation for the line with slope -2/3 and y-intercept (0, 3)

OpenStudy (anonymous):

general form of the equation is y=mx + b here m = -2/3, x = 0 and y = 3 substituting them in general form we get 2 3 = - ---- * 0 + b 3 3 = b so the required equation is 2 y = - ---- x + 3 3